# Describing the Non-Universal Galaxy Merger Timescales in IllustrisTNG: Effects of Host Halo Mass, Baryons, and Sample Selection

### v1

@@ 3.2 Dependence on infall host halo mass
The best-fit parameters and their uncertainties are shown in Figure 4, while the data and best-fit models are presented in Figures 5 and 6, respectively. We find that the parameters indeed depend on $M_{\rm{host}}$ and the results in TNG100-1 and TNG300-1 are now consistent. Specifically, the orbital parameters $\alpha$ and $\gamma$ decrease with increasing $M_{\rm{host}}$, while remaining relatively unchanged for $M_{\rm{host}}<10^{12}M_{\odot}$. This suggests that the dependence of galaxy merger timescales on the initial orbit weakens in higher-mass host halos. For the mass ratio dependence, $\beta$ increases slightly with $M_{\rm{host}}$, from 0.75 to approximately 0.9. Additionally, the scatter $\sigma_{\rm{log}}$ also increases with $M_{\rm{host}}$, further indicating a weaker connection between merger timescales and initial conditions in more massive host halos.
Comparing our results with previous findings from Jiang et al. (2008) and Boylan-Kolchin et al. (2008), we find that for $M_{\rm{host}}<10^{12}M_{\odot}$, the parameters $\alpha$ and $\gamma$ approach the values of 1.9 and 1.0, respectively, as reported in the idealized simulations of Boylan-Kolchin et al. (2008). At higher host masses, our results align with the conclusion of Jiang et al. (2008), which suggests that the dependence on initial orbital conditions becomes relatively weak. This may explain the discrepancies between these two studies as they are studying systems in different conditions. It is reasonable that in the idealized simulations of Boylan-Kolchin et al. (2008) merger timescales show the strongest connection to the initial conditions. In realistic simulations, violent processes in host halos, such as major mergers and interactions between satellite galaxies, can make the infall process more collisional, thereby erasing initial conditions and weakening their connection to merger timescales. According to our results, these processes may occur more frequently in more massive host halos.
@@
and a constant $\beta=\beta_{1}$, where $\rm{erf}$ is the error function and $\rm{erfc}$ is the complementary error function. Therefore, the model contains 9 parameters: $\{A_{\rm{h}},M_{\rm{min}},\sigma_{\log M},A_{\rm{min}},\alpha_{\rm{h}},\alpha_{\rm{min}},\gamma_{\rm{h}},\gamma_{\rm{min}},\beta_{1}\}$. We fit the model directly to the merger timescales of the combined TNG100-1 and TNG300-1 sample. The best-fit model is represented by black solid lines in Figure 4, with the corresponding parameters listed in the first row of Table 1. We find that the best-fit model generally follows the trends observed when modeling different $M_{\rm{host}}$ bins separately. Additionally, we obtain a similar $\sigma_{\rm{log}}$ as in the $M_{\rm{host}}$ bin results, indicating that our model provides a sufficiently accurate description.
In Figure 4, we also present $\sigma_{\rm{log}}$ for different $M_{\rm{host}}$ using the parameters from the full-sample fit shown in Figure 3. The $\sigma_{\rm{log}}$ from the full-sample model is higher than that from the $M_{\rm{host}}$-dependent model, confirming that incorporating $M_{\rm{host}}$ dependence improves the model’s accuracy.
@@ 4.3 One-to-one comparison
The differing impact of baryons on either side of $M_{\rm{host}}\sim 10^{12.5}h^{-1}M_{\odot}$ suggests that baryonic effects can influence merger timescales in multiple ways, either extending or shortening them. The dominant effect may vary depending on $M_{\rm{host}}$, leading to the observed trend. Firstly, the presence of a disk structure in the central galaxy can more effectively disrupt infalling satellites, shortening merger timescales—particularly for those on radial orbits that pass through the central region (Toth & Ostriker, 1992; Velazquez & White, 1999; Helmi et al., 2003). Secondly, the presence of baryons in the satellite galaxy may increase the inner density of the subhalo, making it more resistant to tidal disruption. If the tidal forces in the inner region of the host halo are strong enough to disrupt the subhalos in DMO simulations, the survival of the corresponding subhalo in baryonic simulations may lead to longer merger timescales. However, if the tidal forces are not strong enough to completely disrupt the subhalo but instead significantly reduce its mass in the DMO simulation, longer merger timescales in DMO runs are expected, as reported in Boylan-Kolchin et al. (2008). Finally, in halos hosting high-mass elliptical galaxies, strong feedback from the supermassive black hole reduces the matter density in the central region of the host halos (Xu et al., 2024, see Figure 17), primarily within $r<100h^{-1}{\rm{kpc}}$. This reduction in density may weaken the deceleration caused by dynamical friction, as described by Chandrasekhar’s formula, leading to an increase in merger timescales.
@@
Figure 12: Comparison of merger timescales starting at any radius between simulation results and model predictions using best-fit parameters from those starting at $r_{\rm{vir}}$. The logarithmic difference, $\log_{10}[T_{\rm{mod}}/T_{\rm{sim}}]$, is shown as a function of orbit circularity ($\epsilon$), mass ratio ($M_{\rm sat}/M_{\rm host}(<r)$), orbit energy ($r_{c}/r$), host halo mass at the time of crossing $r_{\rm vir}$ ($M_{{\rm host},0}$), and infall radius ($r$).
@@
Figure 13: The enclosed mass fraction, $M_{\rm host}(<r)/M_{\rm host,0}$, is shown as a function of the infall radius, $r/r_{\rm host,0}$. The red line represents the best-fit relation.
@@
Figure 14: Similar to Figure 12, but showing a comparison between simulation results and predictions from the new model, which combines Equations 7, 8, 10, 11, and 12. For reference, the last panel also presents the best-fit results for the model without corrections from Equation 12 for $T_{\rm dy}$. The red dashed lines indicate scatter values of -0.2 and 0.2.
@@
Figure 15: Distributions of $\log_{10}(M_{\rm{sat}}/M_{\rm{host}})$, $\epsilon$, and $r_{c}/r_{\rm{host}}$ at the time of first crossing $r_{\rm{vir}}$ for our infall samples, combining TNG100-1 and TNG300-1. The distributions are further divided into three sub-samples: merged, disrupted, and still surviving.
@@
We conduct a more detailed analysis by comparing the infall history of matched merger events in DMO and baryonic runs. Figure 11 presents two matched events, one in a low-mass host halo hosting a central disk galaxy and the other in a high-mass host halo with a central elliptical galaxy. These two events have similar initial conditions, with $\epsilon\sim 0.13$ and $M_{\rm{sat}}/M_{\rm{cen}}\sim 0.1$. Notably, the differences between the DMO and baryonic runs exhibit opposite trends in high-mass and low-mass halos. In the low-mass halo, the subhalo in the baryonic run loses a significant amount of mass after its first pericenter passage and merges with the central galaxy immediately after the second passage. In contrast, in the DMO run, the subhalo survives through four pericenter passages. In the high-mass halo, the subhalo in the DMO run experiences a more rapid loss of mass and sinks to the center faster than in the baryonic run.
@@
Figure 16: Comparison of $\Delta T_{\rm merger}$ between simulation results and model predictions using the best-fit model derived from the merger sample started at arbitrary radii (third row of Table 1). The logarithmic difference, $\log_{10}[\Delta T_{\rm mod}/\Delta T_{\rm sim}]$, is shown as a function of orbit circularity ($\epsilon$), mass ratio ($M_{\rm sat}/M_{\rm host}(<r)$), orbit energy ($r_{c}/r$), and host halo mass at the time of crossing $r_{\rm vir}$ ($M_{{\rm host},0}$). Results for the whole infall sample, merger sample, disruption sample, and surviving sample are presented in different rows. The red dashed lines indicate scatter values of -0.2 and 0.2.
Figure 17: Similar to Figure 16, but here the simulation results are compared to the best-fit model derived from fitting the whole infall sample (fourth row of Table 1).
5 Models for Arbitrary Infall Radii and selection effects
In this section, we extend the merger timescale model to include mergers starting at arbitrary infall radii. Using this revised model, we examine selection effects on merger timescale results by comparing the model derived from the merger sample with that from the entire infall sample.
5.1 Extension to arbitrary infall radii
Since dark matter halos exhibit universal density profiles (Navarro et al., 1997), a natural question arises: can our model be extended to merger events initiated at any radius within the halo? Some studies have claimed validation for their models (Simha & Cole, 2017; Poulton et al., 2021). Therefore, we extend our model in Equation 5 to estimate merger timescales for any initial infall radius $r$ as follows:
@@ 5.1 Extension to arbitrary infall radii
To test the model, we use the merger sample from TNG100-1 and TNG300-1. For each event, we treat each intermediate snapshot during its infall history as the starting point of a new infall event and compute the corresponding initial conditions and merger timescales. This process requires particle data from each snapshot to determine $M_{\rm{host}}(<r)$. In Figure 12, we present the logarithmic difference, $\log_{10}(T_{\rm{mod}}/T_{\rm{sim}})$, between simulation results and model predictions from Equation 8. This difference is shown as a function of $\epsilon$, mass ratio $M_{\rm sat}/M_{\rm host}(<r)$, orbit energy $r_{c}/r$, host halo mass $M_{{\rm host},0}$ at the time of crossing $r_{\rm host,0}=r_{\rm vir}$, and infall radius $r$. We find that after entering the halo ($\log_{10}(r/r_{\rm host,0})<-0.3$), our extended model underestimates the merger timescales, suggesting a transition before and after infall. Moreover, our model significantly overestimates the merger timescales for mass ratios $<10^{-2}$, where fewer cases were present in the previous sample used for parameter constraints. This may be reasonable, as subhalos in low mass ratio events experience weaker dynamical friction but are more susceptible to disruption. As a result, this effect may cause events with low mass ratio to deviate from the previous mass dependence.
@@
which is normalized such that $f_{c}(1)=1$, ensuring consistency with our previous results. The effective host halo mass is then defined as $f_{c}M_{\rm host}(<r)$, which is also presented in Figure 13. To better describe the mass ratio dependence, We replace $[M_{\rm{host}}(<r)/M_{\rm{sat}}]^{\beta}$ with a double power law:
@@
By combining Equations 7, 8, 10, 11, and 12, we develop a model to describe merger timescales for mergers starting at arbitrary infall radii. This model includes 11 parameters: $\{A_{\rm{h}},M_{\rm{min}},\sigma_{\log M},A_{\rm{min}},\alpha_{\rm{h}},\alpha_{\rm{min}},\gamma_{\rm{h}},\gamma_{\rm{min}},\beta_{1},\beta_{2},M_{\beta}\}$. When including $r_{\rm t}$ and $\eta$, the total number of parameters increases to 13. We fit this new model to the merger sample starting at arbitrary infall radii. The best-fit parameters are listed in the third row of Table 1, and the residuals as a function of various parameters are shown in Figure 14. Our new model effectively describes the merger timescales for the arbitrary infall radius sample, achieving $\sigma_{\log}=0.195$, with the scatter showing little to no dependence on any parameters. In Figure 14, we also show the scatter for the best-fit model without the correction for $T_{\rm dy}$ in the $r/r_{\rm host,0}$ panel. Including this correction really helps reduce the bias dependence on $r/r_{\rm host,0}$. In the new model, we find a weaker dependence of $\alpha$ and $\gamma$ on $M_{\rm host,0}$ compared to the previous model, with $\alpha_{h}=0.406$ (previously 0.537) and $\gamma_{h}=-0.026$ (previously 0.193). Notably, $\gamma$ now shows almost no dependence. This suggests that $M_{\rm host,0}$ may not be a reliable indicator of system variation for mergers that do not start at $r_{\rm vir}$, and more fundamental parameters might be needed. However, since our model already performs well and $M_{\rm host,0}$ remains the most accessible quantity, we do not explore alternative options.
5.2 The whole infall sample and selection effects

### v4

@@ 3.2 Dependence on infall host halo mass
The best-fit parameters and their uncertainties are shown in Figure 4, while the data and best-fit models are presented in Figures 14 and 15, respectively. We find that the parameters indeed depend on $M_{\rm{host}}$ and the results in TNG100-1 and TNG300-1 are now consistent. Specifically, the orbital parameters $\alpha$ and $\gamma$ decrease with increasing $M_{\rm{host}}$, while remaining relatively unchanged for $M_{\rm{host}}<10^{12}M_{\odot}$. This suggests that the dependence of galaxy merger timescales on the initial orbit weakens in higher-mass host halos. For the mass ratio dependence, $\beta$ increases slightly with $M_{\rm{host}}$, from 0.75 to approximately 0.9. Additionally, the scatter $\sigma_{\rm{log}}$ also increases with $M_{\rm{host}}$, further indicating a weaker connection between merger timescales and initial conditions in more massive host halos.
Comparing our results with previous findings from Jiang et al. (2008) and Boylan-Kolchin et al. (2008), we find that for $M_{\rm{host}}<10^{12}M_{\odot}$, the parameters $\alpha$ and $\gamma$ approach the values of 1.9 and 1.0, respectively, as reported in the idealized simulations of Boylan-Kolchin et al. (2008). At higher host masses, our results align with the conclusion of Jiang et al. (2008), which suggests that the dependence on initial orbital conditions becomes relatively weak. This may explain the discrepancies between these two studies as they are studying systems in different conditions. It is reasonable that in the idealized simulations of Boylan-Kolchin et al. (2008) merger timescales show the strongest connection to the initial conditions. In realistic simulations, violent processes in host halos, such as major mergers (Drakos et al., 2019a, b) and interactions between satellite galaxies, can make the infall process more collisional, thereby erasing initial conditions and weakening their connection to merger timescales. For example, a sudden shift in the halo center position following a major merger can significantly weaken the connection, as it alters the reference frame used to calculate the orbital parameters. . According to our results, these processes may occur more frequently in more massive host halos.
@@
and a constant $\beta=\beta_{1}$, where $\rm{erf}$ is the error function and $\rm{erfc}$ is the complementary error function. These functional forms based on the error function are chosen to interpolate between two limiting cases: the strongest orbital dependence, as found in idealized simulation by Boylan-Kolchin et al. (2008) with $\alpha=1.9$ and $\gamma=1.0$, and the weakest case of no dependence, corresponding to $\alpha=0$ and $\gamma=0$. Therefore, the model contains 9 parameters: $\{A_{\rm{h}},M_{\rm{min}},\sigma_{\log M},A_{\rm{min}},\alpha_{\rm{h}},\alpha_{\rm{min}},\gamma_{\rm{h}},\gamma_{\rm{min}},\beta_{1}\}$. We fit the model directly to the merger timescales of the combined TNG100-1 and TNG300-1 sample. The best-fit model is represented by black solid lines in Figure 4, with the corresponding parameters listed in the first row of Table 1. We find that the best-fit model generally follows the trends observed when modeling different $M_{\rm{host}}$ bins separately. Additionally, we obtain a similar $\sigma_{\rm{log}}$ as in the $M_{\rm{host}}$ bin results, indicating that our model provides a sufficiently accurate description.
In Figure 4, we also present $\sigma_{\rm{log}}$ for different $M_{\rm{host}}$ using the parameters from the full-sample fit shown in Figure 3. The $\sigma_{\rm{log}}$ from the full-sample model is higher than that from the $M_{\rm{host}}$-dependent model, confirming that incorporating $M_{\rm{host}}$ dependence improves the model’s accuracy. We also note that TNG300-1 shows a smaller $\sigma_{\rm{log}}$ than TNG100-1 at fixed $M_{\rm host}$. This is because the merger samples in TNG100-1 span a broader mass ratio range—extending to lower values—due to a lower $M_{*}^{\rm sat}(z_{\rm infall})$ cut. If merger timescales retain some dependence on $M_{\rm sat}$, this broader range would naturally lead to a larger $\sigma_{\rm{log}}$ in TNG100-1, which appears to be the case. We do not pursue a detailed investigation of the $M_{\rm sat}$ dependence, as the current model already yields a relatively low $\sigma_{\rm{log}}$.
@@ 4.3 One-to-one comparison
The differing impact of baryons on either side of $M_{\rm{host}}\sim 10^{12.5}h^{-1}M_{\odot}$ suggests that baryonic effects can influence merger timescales in multiple ways, either extending or shortening them. The dominant effect may vary depending on $M_{\rm{host}}$, leading to the observed trend. Firstly, the presence of a disk structure in the central galaxy can more effectively disrupt infalling satellites, shortening merger timescales—particularly for those on radial orbits that pass through the central region (Toth & Ostriker, 1992; Velazquez & White, 1999; Helmi et al., 2003; Garrison-Kimmel et al., 2017; Samuel et al., 2020). Secondly, the presence of baryons in the satellite galaxy may increase the inner density of the subhalo, making it more resistant to tidal disruption. If the tidal forces in the inner region of the host halo are strong enough to disrupt the subhalos in DMO simulations, the survival of the corresponding subhalo in baryonic simulations may lead to longer merger timescales. However, if the tidal forces are not strong enough to completely disrupt the subhalo but instead significantly reduce its mass in the DMO simulation, longer merger timescales in DMO runs are expected, as reported in Boylan-Kolchin et al. (2008). Finally, in halos hosting high-mass elliptical galaxies, strong feedback from the supermassive black hole reduces the matter density in the central region of the host halos (Xu et al., 2024, see Figure 17), primarily within $r<100h^{-1}{\rm{kpc}}$. This reduction in density may weaken the deceleration caused by dynamical friction, as described by Chandrasekhar’s formula, leading to an increase in merger timescales.
@@
Figure 8: Comparison of merger timescales starting at any radius between simulation results and model predictions using best-fit parameters from those starting at $r_{\rm{vir}}$. The logarithmic difference, $\log_{10}[T_{\rm{mod}}/T_{\rm{sim}}]$, is shown as a function of orbit circularity ($\epsilon$), mass ratio ($M_{\rm sat}/M_{\rm host}(<r)$), orbit energy ($r_{c}/r$), host halo mass at the time of crossing $r_{\rm vir}$ ($M_{{\rm host},0}$), and infall radius ($r$). The figures are color-coded by the number of arbitrary starting merger events.
@@
Figure 9: The enclosed mass fraction, $M_{\rm host}(<r)/M_{\rm host,0}$, is shown as a function of the infall radius, $r/r_{\rm host,0}$. The red solid and dashed lines represent the best-fit and corrected relation. The figures are color-coded by the number of arbitrary starting merger events.
@@
Figure 10: Similar to Figure 8, but showing a comparison between simulation results and predictions from the new model, which combines Equations 7, 8, 10, 11, and 12. For reference, the last panel also presents the best-fit results for the model without corrections from Equation 12 for $T_{\rm dy}$. The red dashed lines indicate scatter values of -0.2 and 0.2. The figures are color-coded by the number of arbitrary starting merger events.
@@
Figure 11: Distributions of $\log_{10}(M_{\rm{sat}}/M_{\rm{host}})$, $\epsilon$, and $r_{c}/r_{\rm{host}}$ at the time of first crossing $r_{\rm{vir}}$ for our entire infall samples, combining TNG100-1 and TNG300-1. The distributions are further divided into three sub-samples: merged ($j_{\rm{merger}}/j_{\rm{infall}}<0.1$), disrupted ($j_{\rm{merger}}/j_{\rm{infall}}>0.1$, regardless of whether the disruption is physical or artificial), and still surviving at $z=0$.
@@
We conduct a more detailed analysis by comparing the infall history of matched merger events in DMO and baryonic runs. Figure 7 presents two matched events, one in a low-mass host halo hosting a central disk galaxy and the other in a high-mass host halo with a central elliptical galaxy. These two events have similar initial conditions, with $\epsilon\sim 0.13$ and $M_{\rm{sat}}/M_{\rm{cen}}\sim 0.1$. The selected events have relatively radial orbits, which exhibit more pronounced differences between the two runs. The mass ratios are chosen to ensure sufficiently long merger histories for clearer illustration, while avoiding rare cases with very small mass ratios. Notably, the differences between the DMO and baryonic runs exhibit opposite trends in high-mass and low-mass halos. In the low-mass halo, the subhalo in the baryonic run loses a significant amount of mass after its first pericenter passage and merges with the central galaxy immediately after the second passage. In contrast, in the DMO run, the subhalo survives through four pericenter passages. In the high-mass halo, the subhalo in the DMO run experiences a more rapid loss of mass and sinks to the center faster than in the baryonic run.
@@
Figure 12: Comparison of $\Delta T_{\rm merger}$ between simulation results and model predictions using the best-fit model derived from the merger sample started at arbitrary radii (third row of Table 1). The logarithmic difference, $\log_{10}[\Delta T_{\rm mod}/\Delta T_{\rm sim}]$, is shown as a function of orbit circularity ($\epsilon$), mass ratio ($M_{\rm sat}/M_{\rm host}(<r)$), orbit energy ($r_{c}/r$), and host halo mass at the time of crossing $r_{\rm vir}$ ($M_{{\rm host},0}$). Results for the whole infall sample, merger sample, disruption sample, and surviving sample are presented in different rows. The red dashed lines indicate scatter values of -0.2 and 0.2. The figures are color-coded by the number of infall events.
Figure 13: Similar to Figure 12, but here the simulation results are compared to the best-fit model derived from fitting the whole infall sample (fourth row of Table 1). The figures are color-coded by the number of infall events.
5 Models for Arbitrary Starting Radii and selection effects
In this section, we extend the merger timescale model to include mergers starting at arbitrary starting radii. Using this revised model, we examine selection effects on merger timescale results by comparing the model derived from the merger sample with that from the entire infall sample.
5.1 Extension to arbitrary starting radii
Since dark matter halos exhibit universal density profiles (Navarro et al., 1997), a natural question arises: can our model be extended to merger events initiated at any radius within the halo, as the inner regions of halos can be considered as new halos with different overdensities (accounted by $T_{\rm dy}$) and concentrations? Some studies have claimed validation for their models (Simha & Cole, 2017; Poulton et al., 2021). Therefore, we extend our model in Equation 5 to estimate merger timescales for any initial starting radius $r$ as follows:
@@ 5.1 Extension to arbitrary starting radii
To test the model, we use the merger sample from TNG100-1 and TNG300-1. For each event, we treat each intermediate snapshot during its infall history as the starting point of a new event and compute the corresponding initial conditions and merger timescales. This process requires particle data from each snapshot to determine $M_{\rm{host}}(<r)$. In Figure 8, we present the logarithmic difference, $\log_{10}(T_{\rm{mod}}/T_{\rm{sim}})$, between simulation results and model predictions from Equation 8. This difference is shown as a function of $\epsilon$, mass ratio $M_{\rm sat}/M_{\rm host}(<r)$, orbit energy $r_{c}/r$, host halo mass $M_{{\rm host},0}$ at the time of crossing $r_{\rm host,0}=r_{\rm vir}$, and starting radius $r$. We find that after entering the halo ($\log_{10}(r/r_{\rm host,0})<-0.3$), our extended model underestimates the merger timescales, suggesting a transition before and after infall. Moreover, our model significantly overestimates the merger timescales for mass ratios $<10^{-2}$, where fewer cases were present in the previous sample used for parameter constraints. This may be reasonable, as subhalos in low mass ratio events experience weaker dynamical friction but are more susceptible to disruption. As a result, this effect may cause events with low mass ratio to deviate from the previous mass dependence.
@@
which is normalized such that $f_{c}(1)=1$, ensuring consistency with our previous results. The effective host halo mass is then defined as $f_{c}M_{\rm host}(<r)$, which is also presented in Figure 9. To better capture the mass ratio dependence, motivated by the power-law upturn bias seen in Figure 8, we replace $[M_{\rm{host}}(<r)/M_{\rm{sat}}]^{\beta}$ with a double power-law form:
@@
By combining Equations 7, 8, 10, 11, and 12, we develop a model to describe merger timescales for mergers starting at arbitrary starting radii. This model includes 11 parameters: $\{A_{\rm{h}},M_{\rm{min}},\sigma_{\log M},A_{\rm{min}},\alpha_{\rm{h}},\alpha_{\rm{min}},\gamma_{\rm{h}},\gamma_{\rm{min}},\beta_{1},\beta_{2},M_{\beta}\}$. When including $r_{\rm t}$ and $\eta$, the total number of parameters increases to 13. We fit this new model to the merger sample starting at arbitrary starting radii. The best-fit parameters are listed in the third row of Table 1, and the residuals as a function of various parameters are shown in Figure 10. Our new model effectively describes the merger timescales for the arbitrary starting radius sample, achieving $\sigma_{\log}=0.195$, with the scatter showing little to no dependence on any parameters. In Figure 10, we also show the scatter for the best-fit model without the correction for $T_{\rm dy}$ in the $r/r_{\rm host,0}$ panel. Including this correction really helps reduce the bias dependence on $r/r_{\rm host,0}$. In the new model, we find a weaker dependence of $\alpha$ and $\gamma$ on $M_{\rm host,0}$ compared to the previous model, with $\alpha_{h}=0.406$ (previously 0.537) and $\gamma_{h}=-0.026$ (previously 0.193). Notably, $\gamma$ now shows almost no dependence. This suggests that $M_{\rm host,0}$ may not be a reliable indicator of system variation for mergers that do not start at $r_{\rm vir}$, and more fundamental parameters might be needed. However, since our model already performs well and $M_{\rm host,0}$ remains the most accessible quantity, we do not explore alternative options.
5.2 The entire infall sample and selection effects